# Can We Trust AI Benchmarks? An Interdisciplinary Review of Current Issues in AI Evaluation

### v1

@@ 5.4. Narrow Benchmark Diversity and Scope
Previous research has further found that current benchmarking practices suffer from diversity issues, a problem that is also found within the broader AI ecosystem (Gomez et al., 2024). On the one hand, a vast majority of benchmarks focus on text, while other modalities (audio, images, video, and multimodal systems) remain largely unexamined (Rauh et al., 2024; Weidinger et al., 2023; Röttger et al., 2024). This concentration is problematic since AI models are increasingly multimodal in scope. Guldimann et al. (2024) also find that benchmarks addressing user privacy, copyright infringement, and interpretability are currently lacking and incomprehensive, while benchmarks in safety areas dealing with corrigibility and explainability are practically missing all together. This leads the authors to conclude that current practices for evaluating topics like safety and ethics “are often simplistic and brittle, leading to inconclusive results” (Guldimann et al., 2024, p. 3). Koch et al. (2021) further identify a concentration on fewer and fewer benchmark datasets within most task communities and note that dominant benchmarks have been introduced by researchers at just a handful of elite institutions, raising questions about representation diversity in the design of benchmarks. Scholars have also found that current AI safety evaluation practices almost exclusively deal with English content (McIntosh et al., 2024; Röttger et al., 2024), and are frequently based on datasets where minorities are under-represented, despite efforts to diversify them (Simson et al., 2024). This raises concerns regarding the inclusion of multiple perspectives on complex topics like ethics and harm.
@@
Aside from mainly focusing on a small range of tasks (evaluating capabilities of English-speaking language models), previous research has identified that most benchmarks tend to be abstracted out of their social and cultural context (Selbst et al., 2019), and rely on a static, one-time testing logic, where results from single evaluations are taken to represent model capabilities writ large. This has given rise to calls for more multi-layered (Weidinger et al., 2023), longitudinal (Mizrahi et al., 2024), and holistic evaluation methods that make sure that AI models do not just perform well in controlled environments, but also in critical/real world circumstances over time (Ojewale et al., 2024; McIntosh et al., 2024; Chang et al., 2023). In their survey of LLM benchmark inadequacies, McIntosh et al. note that ”current benchmarks generally adopt task-based formats, such as Multiple Choice Questions (MCQs) and dialogue-based evaluations, which tend to be static and do not capture the evolving nature of human-AI interactions” (McIntosh et al., 2024, p. 6). Reuel et al. (2024) further note that failures to re-run evaluations multiple times (using different random seeds and sampling temperatures for example), implies that very little is known about the intra-model variance of benchmarks. As a result, they conclude that it is possible that ”most benchmarks fail to distinguish signal and noise” (Reuel et al., 2024, p. 9). Others have emphasised that AI audits often fail to consider risks associated with multiple (inter)acting AI systems (Birhane et al., 2024), and rarely take human actions and motivations into consideration (Weidinger et al., 2023; Chang et al., 2023; Rauh et al., 2024).
@@
While most benchmarks are designed to tell us something about a model’s success, research has further pointed out that they often reveal little (or nothing) about their particular ways of making mistakes, which is crucial from an AI safety and policy enforcement perspective. As Gehrmann et al. (2023) put it, ”ranking models according to a single quality number is easy and actionable - we simply pick the model at the top of the list - [yet] it is much more important to understand when and why models fail” (Gehrmann et al., 2023, p. 130). For instance, they suggest that a focus on errors and fragilities (as opposed to instances of success) can be useful for developers of smaller models, since ”work on quantifying shortcomings is equally applicable to smaller models and methods that improve model robustness often work on all model sizes” (Gehrmann et al., 2023, p. 131). In this sense, failure-focused benchmarks could play an important role in equalling out the playing field in AI development.
5.5. Economic, Competitive, and Commercial Roots
@@
Another contextual element that has been singled out as important is the competitive and commercial roots of benchmark tests. Previous research has emphasised that capability-oriented benchmarks are deeply embedded in corporate marketing strategies and play an important role in increasing the AI hype, attracting customers and investors, and showcasing how models outperform competitors (Orr and Kang, 2024; Grill, 2024; Zhijia, 2024). As Orr and Kang (2024, p. 1881) put it, benchmarks ”serve as the technological spectacle through which companies such as OpenAI and Google can market their technologies”. Many benchmarks also have origins from within the industry and are capability-oriented and centred around tasks with a high potential economic reward, as opposed to focusing on other goals such as ethics and safety (Ren et al., 2024; Ethayarajh and Jurafsky, 2021).
@@
Previous research has noted that this competitive and corporate embedding discourages thorough self-critique since there is a direct ”incentive mismatch between conducting high-quality evaluations and publishing new models or modelling techniques” (Gehrmann et al., 2023, p. 103) and that the field of AI development is ”turning into a giant leaderboard, where publication depends on numbers and little else (such as insight and explanation)” (Church and Hestness, 2019). While a lack of incentives to disclose weaknesses and limitations is a general problem in science (Smith et al., 2022), it has been pointed out that benchmarks have played an especially central role in naturalizing and solidifying a competitive culture in AI research, which is increasingly approached as a ”sport” (Orr and Kang, 2024). As of late, benchmark evaluations have also become increasingly professionalized and transformed into an industry in itself with the rise of platforms like Kaggle and Grand Challenge, who provide organisational and infrastructural support to AI competitions and increasingly function as infrastructures of power in fields like medical imaging (Luitse et al., 2024). The issue of optimising for high benchmark scores at the expense of insight and explanation is known as a form of SOTA-chasing (Koch et al., 2021) and sometimes described as the ”benchmark effect” (Stewart, 2023). Previous research (Malevé, 2023) has also described benchmarks as an example of what Stengers (2018) calls ”fast track research” which idolises rapid, cumulative publication, thus producing a ”winners curse” in AI development (Sculley et al., 2018).
@@
Risks associated with the competitive and commercial roots of benchmarks have further been linked to the growing influence of industry in AI research, where private businesses share of the biggest AI models has increased from 11% in 2010 to 96% in 2021 (Ahmed et al., 2023). In such a context, researchers have noted that current benchmarking tasks - which are generally highly data-intensive - are especially well suited to fit AI models that have been developed within the industry, whose access to advanced data infrastructures, computing power, valuable datasets, and skilled researchers now vastly exceed those of academic researchers (Ahmed et al., 2023). This concentration of power could potentially stifle robust AI evaluations and hinder the development of AI models that adhere to other aims and goals than commercial ones. Scholars have also warned that if academic researchers continue to uphold data-intensive benchmark tests as the SOTA, there is a risk that their research will become increasingly dependent on technological infrastructures provided by the industry (Koch and Peterson, 2024).
5.6. Rigging, Gaming, and Measure Becoming Target
@@
A closely related issue concerns how benchmark tests can be tricked and gamed. In areas and modalities where best-practice benchmarks are missing (i.e., practically all modalities, except for text-based benchmark evaluations), researchers have noted that there are strong incentives to ”rig” benchmark tests. For instance, Dehghani et al. (2021) survey how know-how and recipes for how to score high on benchmark setups are often widely circulated online. Recently, language models have also been found to be optimised for answering the multiple choice questions that are often part of benchmarks (Alzahrani et al., 2024), and to (either intentionally or unintentionally) ”fake” alignment with ethics or safety goals (Greenblatt et al., 2024). The issue point towards what is known as Goodhart’s law, i.e. the recognition that ”when a measure becomes a target, it ceases to be a good measure” (Strathern, 1997).
@@
One reason why gaming can proceed is that users of benchmarks rarely provide the resources needed to validate and replicate their test results (Bartz-Beielstein et al., 2020; Dehghani et al., 2021; Biderman et al., 2024; Reuel et al., 2024). To examine the relevance and validity of benchmark scores, researchers have emphasised that it is necessary to access information concerning all aspects of the evaluation procedure (such as the original evaluation code and all details concerning the experimental setup) (Biderman et al., 2024). Providing such documentation is far from standard, however, especially when proprietary AI models are concerned. This makes it possible to tweak and cherry-pick benchmark results - a problem that is especially pressing given that subtle ”variations in prompts, formatting or other implementation details can significantly impact the performance and validity of evaluations” (Biderman et al., 2024, p. 3). In their in-depth analysis of 24 SOTA language model benchmarks, Reuel et al. (2024) found that only four provided scripts to replicate the results and that no more than ten performed multiple evaluations or reported the statistical significance of their results.
@@
An increasingly well discussed issue also concerns the problem of ”data contamination” i.e., the risk that the models have either intentionally or unintentionally ingested benchmark datasets during training, which severely questions the integrity of AI tests (Xu et al., 2024; Zhang et al., 2024; Besen, 2024; Magar and Schwartz, 2022; Roberts et al., 2023). The problem - which can be an instance of data leakage (Kaufman et al., 2012) or train-test-overlap (Lewis et al., 2021) - has been known for long, and produces similar effects to those of overfitting and memorization (Tirumala et al., 2022; Magar and Schwartz, 2022), leading to models with low generalization power that perform well on familiar tasks (in-distribution) but fail other tasks with a similar difficulty and distribution shift (out-of-distribution) (Yuan et al., 2023; Xu et al., 2024; Zhang et al., 2024; Besen, 2024; Magar and Schwartz, 2022; Roberts et al., 2023; Narayanan and Kapoor, 2023b). When testing GPT4 on benchmark problems from Codeforces (a website hosting coding competitions) in 2023, for instance, Narayanan and Kapoor (2023b) found that the AI model could regularly solve benchmark problems classified as easy - as long as the problems had been added before 5th September 2021. For problems added later, GPT4 could not get a single question right, suggesting that the model had memorised questions and answers. Similar results have also been identified in multiple other models and benchmarks (Xu et al., 2024; Zhang et al., 2024; Besen, 2024; Magar and Schwartz, 2022; Roberts et al., 2023). Despite the fact that issues with data leaks are so well known that strategies have been developed to avoid it [^5] https://livebench.ai. Last seen 22nd January 2025., there is still a widespread lack of reporting of data contamination tendencies during benchmark tests. In a study from October 2024, for example, Zhang et al. (2024) found that out of 30 analysed models, only 9 reported train test overlap.
@@
Another recently identified issue is called ”sandbagging” and involves a ”strategic underperformance on an evaluation” which occurs when an AI developer intentionally understates a models capability, for instance to avoid becoming a target for AI safety regulation. In a 2024 study, for example, Weij et al. (2024, p. 1) prompted frontier models like GPT-4 and Claude 3 Opus and found that they ”selectively underperform on dangerous capability evaluations, while maintaining performance on general (harmless) capability evaluations”. The researchers also found that that it was possible to fine-tune and adjust both frontier and smaller models to hide specific capabilities or target specific capability scores. The issue could be generalised to popular benchmarks like the Weapons of Mass Destruction Proxy Benchmark (WMDP), and puts the trustworthiness of benchmark evaluations into question, especially in a regulatory context.
5.7. Dubious Community Vetting and Path Dependencies
@@
A related area of research emphasises how benchmarks become naturalised and reach standard status because of the culture and logic of academic citations (Orr and Kang, 2024). For instance, new benchmarks are commonly introduced together with new or updated AI models. If the AI model becomes popular, the benchmark may become widely cited and circulated as a secondary effect, even though the developers of the benchmark did not intended or expect it to become standard. In this way, a benchmark that might have been selected for practical purpose or due to subjective preferences may come to have substantial technical and cultural influence as a result of community vetting, even though its suitability as a yardstick for ethics, safety, or performance may be questioned (Orr and Kang, 2024; Orr and Crawford, 2024b). For instance, Denton et al. (2021) show how this was partly the case when the ImageNet dataset - a key reference point in the performance testing of computer vision models - became standard, following the unforeseen success of the ImageNet Large Scale Visual Recognition Challenge hosted by Toronto University in 2010. Likewise, the so-called Lena test image - also central to computer vision benchmark tests - was taken from the centerfold of a November 1972 Playboy magazine and catapulted into computer history since ”someone happened to walk in with a recent issue of Playboy” at a time of need, according to what has been described as the most credible origin story (Mulvin, 2021, p. 80). Since then, the ImageNet dataset and its associated benchmark challenge has been become a symbol for dataset bias (Denton et al., 2021), while the Lena test image has come to serve as a prime example of the role of whiteness and women’s sexualised bodies in the standardisation of digital visual culture (Mulvin, 2021). These examples highlight how benchmarks are fundamentally cultural and political products whose power and influence may and be (uncritically and problematically) reinforced through community vetting.
@@
Aside from bias and representational issues, Schlangen (2020, p. 1) argues that what is typically missing and left implicit in the peer-review-fuelled process of benchmark use is the argumentation for why scoring well on a particular benchmark ”constitutes progress, and progress towards what”. Instead, researchers are expected to routinely demonstrate performance on dominant benchmarks, despite the fact that more task-specific benchmarks may be more technically appropriate (Koch et al., 2021). In this way, peer-washing serves to ”maintain datasets as authoritative proxies even when they are shown to be harmful or problematic” (Orr and Crawford, 2024b, p. 4966). New benchmarks often have a difficult time to gain traction because of the dominance and authority of well-cited benchmarks (Jaton, 2021), and in a study of 3765 benchmarks, Ott et al. (2022) found that only a small minority of the proposed benchmark solutions reached widespread adoption. This problem is also recognised within the industry. For instance, researchers at Google’s Brain Team describe what they call a ”benchmark lottery” which ”postulates that many factors, other than fundamental algorithmic superiority, may lead to a method being perceived as superior” (Dehghani et al., 2021, p. 1). Ironically, researchers have simultaneously found that a majority of influential benchmarks have been released as preprints without going through rigorous academic peer-review (McIntosh et al., 2024, p. 6). In other words, problems with peer-washing both concern the poor quality control that many benchmarks are subjected to, and the self-fulfilling and sometimes excessive ways through which certain benchmarks are propelled into standards.
@@
Bao et al. (2022) further note that most papers that introduce benchmarks have origins in the field of machine learning and are mainly focused on methods. This means that the content of benchmarks datasets are often considered secondary, since the datasets are merely there to provide a baseline for comparison between different evaluation methods. When benchmarks are applied to real and specific use-cases - such as evaluating the fairness of algorithmic risk assessment instruments within the criminal justice system - the downstream effects of such a lack of concern for datasets can have worrying effects. As Bao et al. (2022) put it, a paper on benchmarks can be of high quality in a pure AI/ML methods sense, but irrelevant, dangerous, or harmful when applied in circumstances such as the criminal justice system, since it may introduce and perpetuate harms and mistranslations. The researchers also suggest that the current peer-review system implies that benchmarks that are primarily relevant from a methods/machine learning perspective will be cited far more often than benchmarks that are relevant for use in specific, real-life use-cases. Looking too closely at citation counts when determining the quality and relevance of benchmarks may thus effectively lead users astray and complicate identifying benchmarks with a high practical utility.
@@
Problems with dubious community vetting become especially worrying given that benchmarks create ”path dependencies” in AI research, meaning they reinforce certain methodologies and research goals, while stifling those that do not align with the logic of dominant benchmark tests (Blili-Hamelin and Hancox-Li, 2023). In particular, Koch and Peterson (2024, p. 3) warn against the tendency for benchmarks to favour a form of task-driven scientific monoculture that privileges immediate, explicit, formal, quantitative, and easily-interpretable evaluation mechanisms that ”prioritize one or a few key epistemic values (e.g., accuracy, safety),” at the expense of a broader and more complex vision of scientific progress. The authors note how a series of methodological paradigms in AI research (such as boosting, Bayesian networks, Bayesian non-parametrics, and support vector machines) have rapidly faded away following the boom in deep learning research around the year of 2014, and partly attribute this to the ”epistemic narrowness” of current benchmark tests, which are well-suited for deep learning models (whose performance can be reliably increased through scaling), but less favourable for other methodological paradigms in AI. ”When scientists can only gain high-status publications by demonstrating SOTA accuracy,” writes Koch and Peterson (2024, p. 30), ”the safest research choice becomes incrementally advancing proven methods, not innovating new ones”. According to Koch, the cost of current benchmark practices - which privileges scale, compute, and the use of larger and larger training datasets - for instance includes repeated privacy and copyright violations, emotional harms caused to under-paid data workers, and increasing ecological and environmental pressures due to high energy consumption.
5.8. Rapid AI Development and Benchmark Saturation

### v2

@@ 5.4 Narrow Benchmark Diversity and Scope
Previous research has further found that current benchmarking practices suffer from diversity issues, a problem that is also found within the broader AI ecosystem [Gomez et al., 2024]. On the one hand, a vast majority of benchmarks focus on text, while other modalities (audio, images, video, and multimodal systems) remain largely unexamined [Rauh et al., 2024; Weidinger et al., 2023; Röttger et al., 2024]. This concentration is problematic since AI models are increasingly multimodal in scope. Guldimann et al. [2024] also find that benchmarks addressing user privacy, copyright infringement, and interpretability are currently lacking and incomprehensive, while benchmarks in safety areas dealing with corrigibility and explainability are practically missing all together. This leads the authors to conclude that current practices for evaluating topics like safety and ethics “are often simplistic and brittle, leading to inconclusive results” [Guldimann et al., 2024, p. 3]. Koch et al. [2021] further identify a concentration on fewer and fewer benchmark datasets within most task communities and note that dominant benchmarks have been introduced by researchers at just a handful of elite institutions, raising questions about representation diversity in the design of benchmarks. Scholars have also found that current AI safety evaluation practices almost exclusively deal with English content [McIntosh et al., 2024; Röttger et al., 2024; Poelman and Lhoneux, 2024], and are frequently based on datasets where minorities are under-represented, despite efforts to diversify them [Simson et al., 2024]. This raises concerns regarding the inclusion of multiple perspectives on complex topics like ethics and harm.
@@
Aside from mainly focusing on a small range of tasks, previous research has identified that most benchmarks tend to be abstracted out of their social and cultural context [Selbst et al., 2019; Lum et al., 2024], are aggregated in problematic ways [Burnell et al., 2023], and rely on a static, one-time testing logic, where results from single evaluations are taken to represent model capabilities writ large. This has given rise to calls for more multi-layered [Weidinger et al., 2023], longitudinal [Mizrahi et al., 2024], and holistic evaluation methods that can be reproduced [Kapoor et al., 2024], and prove that AI models do not just perform well in controlled environments, but also in critical/real-world circumstances over time [Ojewale et al., 2024; McIntosh et al., 2024; Chang et al., 2023]. In their survey of LLM benchmark inadequacies, McIntosh et al. note that "current benchmarks generally adopt task-based formats, such as Multiple Choice Questions (MCQs) and dialogue-based evaluations, which tend to be static and do not capture the evolving nature of human-AI interactions" [McIntosh et al., 2024, p. 6]. Reuel et al. [2024] further note that failures to re-run evaluations multiple times (using different random seeds and sampling temperatures, for example) imply that very little is known about the intra-model variance of benchmarks. As a result, they conclude that it is possible that "most benchmarks fail to distinguish signal and noise" [Reuel et al., 2024, p. 9]. Others have emphasised that AI audits often fail to consider risks associated with multiple (inter)acting AI systems [Birhane et al., 2024], and rarely take human actions and motivations into consideration [Weidinger et al., 2023; Chang et al., 2023; Rauh et al., 2024].
@@
While most benchmarks are designed to tell us something about a model’s success, research has further pointed out that they often reveal little (or nothing) about their particular ways of making mistakes, which is crucial from an AI safety and policy enforcement perspective. As Gehrmann et al. [2023] put it, "ranking models according to a single quality number is easy and actionable - we simply pick the model at the top of the list - [yet] it is much more important to understand when and why models fail" [Gehrmann et al., 2023, p. 130]. For instance, they suggest that a focus on errors and fragilities (as opposed to instances of success) can be useful for developers of smaller models, since "work on quantifying shortcomings is equally applicable to smaller models and methods that improve model robustness often work on all model sizes" [Gehrmann et al., 2023, p. 131]. In this sense, failure-focused benchmarks could play an important role in equalling out the playing field in AI development.
5.5 Economic, Competitive, and Commercial Roots
@@
Another contextual element that has been singled out as important is the competitive and commercial roots of benchmark tests. Previous research has emphasised that capability-oriented benchmarks are deeply embedded in corporate marketing strategies and play an important role in increasing the AI hype, attracting customers and investors, and showcasing how models outperform competitors [Orr and Kang, 2024; Grill, 2024; Zhijia, 2024]. As Orr and Kang [2024, p. 1881] put it, benchmarks "serve as the technological spectacle through which companies such as OpenAI and Google can market their technologies". Many benchmarks also have origins from within the industry and are capability-oriented and centred around tasks with a high potential economic reward, as opposed to focusing on other goals such as ethics and safety [Ren et al., 2024; Ethayarajh and Jurafsky, 2021].
@@
Previous research has noted that this competitive and corporate embedding discourages thorough self-critique since there is a direct "incentive mismatch between conducting high-quality evaluations and publishing new models or modelling techniques" [Gehrmann et al., 2023, p. 103] and that the field of AI development is "turning into a giant leaderboard, where publication depends on numbers and little else (such as insight and explanation)" [Church and Hestness, 2019]. While a lack of incentives to disclose weaknesses and limitations is a general problem in science [Smith et al., 2022], it has been pointed out that benchmarks have played an especially central role in naturalizing and solidifying a competitive culture in AI research, which is increasingly approached as a "sport" [Orr and Kang, 2024]. As of late, benchmark evaluations have also become increasingly professionalized and transformed into an industry in itself with the rise of platforms like Kaggle and Grand Challenge, who provide organisational and infrastructural support to AI competitions and increasingly function as infrastructures of power in fields like medical imaging [Luitse et al., 2024]. The issue of optimising for high benchmark scores at the expense of insight and explanation is known as a form of SOTA-chasing [Koch et al., 2021] and sometimes described as the "benchmark effect" [Stewart, 2023]. Previous research has also described benchmarking as an example of what Stengers [2018] calls "fast track research" which idolises rapid, cumulative publication [Malevé, 2023], thus producing a "winners curse" in AI development [Sculley et al., 2018].
@@
Risks associated with the competitive and commercial roots of benchmarks have further been linked to the growing influence of industry in AI research, where private businesses share of the biggest AI models has increased from 11% in 2010 to 96% in 2021 [Ahmed et al., 2023]. In such a context, researchers have noted that current benchmarking tasks - which are generally highly data-intensive - are especially well suited to fit AI models that have been developed within the industry, whose access to advanced data infrastructures, computing power, valuable datasets, and skilled researchers now vastly exceed those of academic researchers [Ahmed et al., 2023]. This concentration of power could potentially stifle robust AI evaluations and hinder the development of AI models that adhere to other aims and goals than commercial ones. Scholars have also warned that if academic researchers continue to uphold data-intensive benchmark tests as the SOTA, there is a risk that their research will become increasingly dependent on technological infrastructures provided by the industry [Koch and Peterson, 2024].
5.6 Rigging, Gaming, and Measure Becoming Target
@@
A closely related issue concerns how benchmark tests can be tricked and gamed. In areas and modalities where best-practice benchmarks are missing (i.e., practically all modalities, except for text-based benchmark evaluations), researchers have noted that there are strong incentives to "rig" benchmark tests. For instance, Dehghani et al. [2021] survey how know-how and recipes for how to score high on benchmark setups are often widely circulated online. Recently, language models have also been found to be optimised for answering the multiple choice questions that are often part of benchmarks [Alzahrani et al., 2024], and to (either intentionally or unintentionally) "fake" alignment with ethics or safety goals [Greenblatt et al., 2024] and hide their true capabilities and objectives - also known as scheming [Meinke et al., 2024]. The issue points towards what is known as Goodhart’s law: “when a measure becomes a target, it ceases to be a good measure” [Strathern, 1997].
@@
One reason why gaming can proceed is that users of benchmarks rarely provide the resources needed to validate and replicate their test results [Bartz-Beielstein et al., 2020; Dehghani et al., 2021; Biderman et al., 2024; Reuel et al., 2024] and that they rely on black-box (as opposed to white-box) access to AI models, which undercuts rigorous audits [Casper et al., 2024]. To examine the relevance and validity of benchmark scores, researchers have further emphasised that it is necessary to access information concerning all aspects of the evaluation procedure, such as the original evaluation code and all details concerning the experimental setup) [Biderman et al., 2024]. Providing such documentation is far from standard, however, especially when proprietary AI models are concerned. This makes it possible to tweak and cherry-pick benchmark results - a problem that is especially pressing given that subtle "variations in prompts, formatting or other implementation details can significantly impact the performance and validity of evaluations" [Biderman et al., 2024, p. 3]. In their in-depth analysis of 24 SOTA language model benchmarks, Reuel et al. [2024] found that only four provided scripts to replicate the results and that no more than ten performed multiple evaluations or reported the statistical significance of their results.
@@
An increasingly well discussed issue also concerns the problem of "data contamination" i.e., the risk that the models have either intentionally or unintentionally ingested benchmark datasets during training, which severely questions the integrity of AI tests [Xu et al., 2024a; Zhang et al., 2024; Besen, 2024; Magar and Schwartz, 2022; Roberts et al., 2023; Yang et al., 2023]. The problem - which can be an instance of data leakage [Kaufman et al., 2012; Xu et al., 2024b] or train-test-overlap [Lewis et al., 2021] - has been known for long, and produces similar effects to those of overfitting and memorization [Tirumala et al., 2022; Magar and Schwartz, 2022], leading to models with low generalization power that perform well on familiar tasks (in-distribution) but fail other tasks with a similar difficulty and distribution shift (out-of-distribution) [Yuan et al., 2023; Xu et al., 2024a; Zhang et al., 2024; Besen, 2024; Magar and Schwartz, 2022; Roberts et al., 2023; Narayanan and Kapoor, 2023a]. When testing GPT4 on benchmark problems from Codeforces (a website hosting coding competitions) in 2023, for instance, Narayanan and Kapoor [2023a] found that the AI model could regularly solve benchmark problems classified as easy - as long as the problems had been added before 5th September 2021. For problems added later, GPT4 could not get a single question right, suggesting that the model had memorised questions and answers. Similar results have also been identified in multiple other models and benchmarks [Xu et al., 2024a; Zhang et al., 2024; Besen, 2024; Magar and Schwartz, 2022; Roberts et al., 2023]. Despite the fact that issues with data leaks are so well known that strategies have been developed to avoid it [^5] https://livebench.ai. Last seen May 25, 2025., there is still a widespread lack of reporting of data contamination tendencies during benchmark tests. In a study from October 2024, for example, Zhang et al. [2024] found that out of 30 analysed models, only 9 reported train test overlap.
@@
Another recently identified issue is called "sandbagging" and involves a "strategic underperformance on an evaluation" which occurs when an AI developer intentionally understates a models capability, for instance to avoid becoming a target for AI safety regulation. In a 2024 study, for example, Weij et al. [2024, p. 1] prompted frontier models like GPT-4 and Claude 3 Opus and found that they "selectively underperform on dangerous capability evaluations, while maintaining performance on general (harmless) capability evaluations". The researchers also found that that it was possible to fine-tune and adjust both frontier and smaller models to hide specific capabilities or target specific capability scores. The issue could be generalised to popular benchmarks like the Weapons of Mass Destruction Proxy Benchmark (WMDP), and puts the trustworthiness of benchmark evaluations into question, especially in a regulatory context.
5.7 Dubious Community Vetting and Path Dependencies
@@
A related area of research emphasises how benchmarks become naturalised and reach standard status because of the culture and logic of academic citations [Orr and Kang, 2024]. For instance, new benchmarks are commonly introduced together with new or updated AI models. If the AI model becomes popular, the benchmark may become widely cited and circulated as a secondary effect, even though the developers of the benchmark did not intended or expect it to become standard. In this way, a benchmark that might have been selected for practical purpose or due to subjective preferences may come to have substantial technical and cultural influence, even though its suitability as a yardstick for ethics, safety, or performance may be questioned [Orr and Kang, 2024; Orr and Crawford, 2024a]. For instance, Denton et al. [2021] show how this was partly the case when the ImageNet dataset - a key reference point in the performance testing of computer vision models - became standard, following the unforeseen success of the ImageNet Large Scale Visual Recognition Challenge hosted by Toronto University in 2010. Likewise, the so-called Lena test image - also central to computer vision benchmark tests - was taken from the centerfold of a November 1972 Playboy magazine and catapulted into computer history since "someone happened to walk in with a recent issue of Playboy" at a time of need, according to what has been described as the most credible origin story [Mulvin, 2021, p. 80]. Since then, the ImageNet dataset and its associated benchmark challenge has been become a symbol for dataset bias [Denton et al., 2021], while the Lena test image has come to serve as a prime example of the role of whiteness and women’s sexualised bodies in the standardisation of digital visual culture [Mulvin, 2021]. These examples highlight how benchmarks are fundamentally cultural and political products whose power and influence may and be (uncritically and problematically) reinforced through community vetting.
@@
Aside from bias and representational issues, Schlangen [2020, p. 1] argues that what is typically missing and left implicit in the peer-review-fuelled process of benchmark use is the argumentation for why scoring well on a particular benchmark "constitutes progress, and progress towards what". Instead, researchers are expected to routinely demonstrate performance on dominant benchmarks, despite the fact that more task-specific benchmarks may be more technically appropriate [Koch et al., 2021]. In this way, peer-washing serves to "maintain datasets as authoritative proxies even when they are shown to be harmful or problematic" [Orr and Crawford, 2024a, p. 4966]. New benchmarks often have a difficult time to gain traction because of the dominance and authority of well-cited benchmarks [Jaton, 2021], and in a study of 3765 benchmarks, Ott et al. [2022] found that only a small minority of the proposed benchmark solutions reached widespread adoption. This problem is also recognised within the industry. For instance, researchers at Google’s Brain Team describe what they call a "benchmark lottery" which "postulates that many factors, other than fundamental algorithmic superiority, may lead to a method being perceived as superior" [Dehghani et al., 2021, p. 1]. Ironically, researchers have simultaneously found that a majority of influential benchmarks have been released as preprints without going through rigorous academic peer-review [McIntosh et al., 2024, p. 6]. In other words, problems with peer-washing both concern the poor quality control that many benchmarks are subjected to, and the self-fulfilling and sometimes excessive ways through which certain benchmarks are propelled into standards.
@@
Bao et al. [2022] further note that most papers that introduce benchmarks have origins in the field of machine learning and are mainly focused on methods. This means that the content of benchmarks datasets are often considered secondary, since the datasets are merely there to provide a baseline for comparison between different evaluation methods. When benchmarks are applied to real and specific use-cases - such as evaluating the fairness of algorithmic risk assessment instruments within the criminal justice system - the downstream effects of such a lack of concern for datasets can have worrying effects. As Bao et al. [2022] put it, a paper on benchmarks can be of high quality in a pure AI/ML methods sense, but irrelevant, dangerous, or harmful when applied in circumstances such as the criminal justice system, since it may introduce and perpetuate harms and mistranslations. The researchers also suggest that the current peer-review system implies that benchmarks that are primarily relevant from a methods/machine learning perspective will be cited far more often than benchmarks that are relevant for use in specific, real-life use-cases. Looking too closely at citation counts when determining the quality and relevance of benchmarks may thus effectively lead users astray and complicate identifying benchmarks with a high practical utility.
@@
Problems with dubious community vetting become especially worrying given that benchmarks create "path dependencies" in AI research, meaning they reinforce certain methodologies and research goals, while stifling those that do not align with the logic of dominant benchmark tests [Blili-Hamelin and Hancox-Li, 2023]. In particular, Koch and Peterson [2024, p. 3] warn against the tendency for benchmarks to favour a form of task-driven scientific monoculture that privileges immediate, explicit, formal, quantitative, and easily-interpretable evaluation mechanisms that "prioritize one or a few key epistemic values (e.g., accuracy, safety)," at the expense of a broader and more complex vision of scientific progress. The authors note how a series of methodological paradigms in AI research (such as boosting, Bayesian networks, Bayesian non-parametrics, and support vector machines) have rapidly faded away following the boom in deep learning research around the year of 2014, and partly attribute this to the "epistemic narrowness" of current benchmark tests, which are well-suited for deep learning models (whose performance can be reliably increased through scaling), but less favourable for other methodological paradigms in AI. "When scientists can only gain high-status publications by demonstrating SOTA accuracy," writes Koch and Peterson [2024, p. 30], "the safest research choice becomes incrementally advancing proven methods, not innovating new ones". According to Koch, the cost of current benchmark practices - which privileges scale, compute, and the use of larger and larger training datasets - for instance includes repeated privacy and copyright violations, emotional harms caused to under-paid data workers, and increasing ecological and environmental pressures due to high energy consumption.
5.8 Rapid AI Development and Benchmark Saturation